For a converging lens with two curved surfaces, the radius of curvature for both surfaces is 13.1 cm. If the focal length is 16.5 cm, what must the index of refraction be?
Answer #1

Solution) R1 = R2 = 13.1 cm

f = 16.5 cm

n = ?

(1/f) = ( n - 1 )((1/R1) + (1/R2))

(1/16.5) = ( n - 1 )((1/13.1) + (1/13.1))

(1/16.5) = ( n - 1 )(2/13.1)

n - 1 = (13.1)/(2×16.5)

n - 1 = 0.396

n = 1 + 0.396

n = 1.396